I. Seven letter words are formed using the letters of the word BLACKER.
(a)How many which end in R? (b)How many in which E immediately follows K?
(c)How many do not begin with B?(d)How many in which the vowels are separated by exactly two letters? (e)How many in which the vowels are seperated by two or more letters?
I really appreciate any help

Found 2 solutions by edjones, sudhanshu_kmr:
Answer by edjones(8007)

(a)How many which end in R?
There are 7 letters in the word and one R which is at the end.
6!*1=720
.
Ed

Answer by sudhanshu_kmr(1152)

a)
no. of words end in R = 6! = 720

b)
no. of words in which E immediately follows k = 6! = 720
(assume KE as a single letter, now total 6 letters )

c)
total no. of words with 7 letters = 7!

no. of words begin with B = 6! = 720

so, no. of words not begin with B = 7! - 6! = 4320


d)
here vowels must be in either of these positions.......
(1 and 4),(2 and 5), (3 and 6) or (4 and 7)

no. of ways to arrange vowels = 4*2 = 8 (2 ways for each position)
no. of ways to arrange 5 consonants = 5!
total no. of ways = 120 *8 = 960


e)
try to understand the concept and try to solve it yourself..
you are welcome to contact for any assistance...
sudhanshu.cochin@gmail.com
best of luck........